# Estimation of relative risk, odds ratio and their logarithms with guaranteed accuracy and controlled sample size ratio

### v1

@@ A.1 For relative risk and log relative risk
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\lambda\theta\left(\frac{1}{A}+\mu_{2}+\xi\right)+\frac{1}{A}+r_{1}+\mu_{1}+\xi,$ (173)
@@
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\phi^{2}l_{1}l_{2}$ $\approx\frac{1}{\lambda\theta}\left(\frac{1}{A}+r_{1}+\mu_{1}+\xi\right)^{\!2}\frac{2r_{1}+\mu_{1}-\mu_{2}-1}{r_{1}(r_{1}+\mu_{1}-\mu_{2}-2)}.$ (174)
In practice, the target $A$ will typically be smaller, or much smaller, than $1$. For example, for RR a relative RMSE of $20\%$ corresponds to $A=0.04$. On the other hand, $\xi$ equals $1$, and the values of $r_{1}$ obtained from (81) are small compared to $1/A$, as seen in Figure 3. In the example, $A=0.04$ gives $r_{1}=5$. Thus, from (167), (168), (171)–(174) the following behavior is observed, both for RR ($\mu_{1}=2$, $\mu_{1}=0$) and for LRR ($\mu_{1}=\mu_{2}=1$): if $\lambda\theta\approx 1$ the four terms $\operatorname{Var}[M_{1}/l_{1}]$, $\operatorname{Var}[M_{2}/l_{2}]$, $\operatorname{E}[\operatorname{Var}[N_{1}/l_{1}\mid s_{1}]]$ and $\operatorname{E}[\operatorname{Var}[N_{2}/l_{2}\mid s_{2}]]$ are significantly smaller than $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$ and $\operatorname{Var}[\operatorname{E}[N_{2}/l_{2}\mid s_{2}]]$; if $\lambda\theta>1$ the term $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$ is much greater than the other five; and if $\lambda\theta<1$ the term $\operatorname{Var}[\operatorname{E}[N_{2}/l_{2}\mid s_{2}]]$ is much greater than the others. The conclusion is that, regardless of $\lambda\theta$, one or both of the terms $\operatorname{Var}[\operatorname{E}[N_{i}/l_{i}\mid s_{i}]]$, $i=1,2$ in (166) dominate. This means that to account for the variability of $\Delta$ it is sufficient, with small approximation error, to only consider the variability of $s_{1}$, $s_{2}$, and disregard that of $M_{1}$, $M_{2}$ and that of $N_{1}$, $N_{2}$ conditioned on $s_{1}$, $s_{2}$. Therefore, to compute $\operatorname{E}[|\Delta|]$, the variables $M_{1}$, $M_{2}$, $N_{1}$, $N_{2}$ in (95) can be replaced by their conditional means given $s_{1}$, $s_{2}$, which yields (97).
A.2 For odds ratio and log odds ratio
The decomposition (166) obtained for $\operatorname{Var}[\Delta]$ in RR and LRR also applies in OR and LOR, but the expressions for the six summands are different. To compute $\operatorname{Var}[M_{i}/l_{i}]$, $i=1,2$ the law of total variance [5, theorem 12.2.6] is used:
@@ A.2 For odds ratio and log odds ratio
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Subsection 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, which yields
@@
Thus, for a given $\bar{M}_{i}$, the average number of required inputs is
@@
The term $\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]$ in (175) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and using (176),
@@
$\operatorname{Var}\left[M_{i}\right]=\frac{(14\bar{p}_{i}^{2}-23\bar{p}_{i}+9)r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}=\frac{(9-14\bar{p}_{i})r_{i}}{4\bar{p}_{i}^{2}},$ (180)
and therefore, taking into account (143),
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}\right]\bar{\phi}^{2}l_{1}l_{2}$ $=\frac{(9-14\bar{p}_{1})r_{1}}{4\lambda\bar{\theta}},$ (181)
$\operatorname{Var}\left[\frac{M_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}$ $=\frac{\lambda\bar{\theta}(9-14\bar{p}_{2})r_{1}}{4}.$ (182)
For the other four summands in (166) it is helpful, as in Appendix A.1, to approximate the variances of $X$ and $1/X$ by their values for small $\bar{\phi}$, which in this case yields
$\operatorname{Var}[X]$ $\approx\frac{r_{2}(r_{1}+r_{2}-1)\bar{\theta}^{2}}{(r_{1}-2)(r_{1}-1)^{2}},$ (183)
$\operatorname{Var}[1/X]$ $\approx\frac{r_{1}(r_{1}+r_{2}-1)}{(r_{2}-2)(r_{2}-1)^{2}\bar{\theta}^{2}}.$ (184)
The term $\operatorname{E}[\operatorname{Var}[N_{1}/l_{1}\mid X]]$ is obtained, making use of (13) again, as
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{\operatorname{E}[\operatorname{Var}[N_{1}^{\prime}\mid s_{1}]]\bar{\phi}^{2}}{\lambda}+\frac{\operatorname{E}[\operatorname{Var}[N_{1}^{\prime\prime}\mid s_{1}]]\bar{\phi}^{2}}{\lambda}\\ &\quad=\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[(s_{1}-\alpha)]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}\\ &=\frac{(p_{1}^{3}+(1-p_{1})^{3})\operatorname{E}[s_{1}]\bar{\phi}^{2}-\alpha p_{1}^{3}\bar{\phi}^{2}}{\lambda\bar{p}_{1}^{2}}=\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}},\end{split}$ (185)
with $\alpha$ equal to $2$ for OR and 00 for LOR. Using (37), including the rounding term $\xi$, and then substituting (39), (40), (136) and (142)–(145),
$\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}\mathrel{}\middle|\mathrel{}s_{1}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\frac{(a_{1}+\xi+b_{1}\operatorname{E}[X])(1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}\\ \approx\left(\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(1-3\bar{p}_{1})-\frac{\alpha p_{1}^{3}}{\lambda\bar{\theta}}.$ (186)
As for $\operatorname{Var}[\operatorname{E}[N_{1}/l_{1}\mid s_{1}]]$, from (11), (37), (40), (142)–(145), and (183),
[…]
Analogously, using (11), (13), (38), (41), (42), (137), (142)–(145) and (184),
@@
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\lambda\bar{\theta}\left((a_{2}+\xi+b_{2}\operatorname{E}[1/X])(1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\quad\approx\left(\lambda\bar{\theta}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)(1-3\bar{p}_{2})-\alpha\lambda\bar{\theta}(1-p_{2})^{3},\end{split}$ (188)
@@
$\begin{split}&\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\approx\lambda\bar{\theta}b_{2}^{2}\operatorname{Var}[1/X]\approx\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}\,\frac{2r_{1}-1}{r_{1}(r_{1}-2)}.\end{split}$ (189)
From (181), (182), (186)–(189), a similar assessment as in Appendix A.1 can be made, and the conclusion is the same: for typical values of $A$, both for OR ($\mu_{1}=\mu_{2}=2$, $\alpha=2$) and LOR ($\mu_{1}=\mu_{2}=5/4$, $\alpha=0$), and for any $\lambda$ and $\bar{\theta}$, one or both of the terms $\operatorname{Var}[\operatorname{E}[N_{i}/l_{i}\mid s_{i}]]$ in (166) dominate the sum, which implies that (153) is a good approximation for small $\bar{\phi}$.

### v2

@@ B.2 For odds ratio and log odds ratio
$\begin{split}\displaystyle\operatorname{Var}\left[\operatorname{E}\left[\Delta\mid\bar{M}_{1},\bar{M}_{2}\right]\right]&=\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right].\end{split}$ (175)
On the other hand, making use of the law of total variance again,
$\operatorname{E}\left[\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right].$ (176)
Combining (174)–(176) with (173), and noting that conditioning $N_{1}$ or $N_{2}$ on $\bar{M}_{1}$, $\bar{M}_{2}$ is equivalent to conditioning on $s_{1}$, $s_{2}$, the following expression is obtained for $\operatorname{Var}[\Delta]$:
@@
$\begin{split}\displaystyle\operatorname{Var}\left[\Delta\right]&=\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]+\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]+\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\\ &\quad+2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right].\end{split}$ (177)
This is similar to the decomposition (165) in RR and LRR, where only the covariance term is different. The approximate expressions of the summands for small $\phi$ are also different from that case, as will be computed next. Then, as in RR an LRR, it will be shown that for OR and LOR the third summand dominates the other three.
The first summand in (177) is written, thanks to the independence of $M_{1}$ and $M_{2}$, as
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]=\frac{\operatorname{Var}[M_{1}]}{l_{1}^{2}}+\frac{\operatorname{Var}[M_{2}]}{l_{2}^{2}}.$ (178)
From the law of total variance, $\operatorname{Var}[M_{i}]$, $i=1,2$ is expressed as
@@
The first stage applies IBS to samples with parameter $\bar{p}_{i}$, using a number $\bar{M}_{i}$ of those samples, of which $r_{i}$ are successes. Each sample with parameter $\bar{p}_{i}$ is generated by the Bernoulli factory described in Section 4.1, taking samples with parameter $p_{i}$ as inputs. With this factory, the average number of inputs needed to produce an output is $3/2$; a success output always uses $2$ inputs, and a failure output uses either $1$ or $2$ inputs. Let $\pi_{i}$ denote the probability that $2$ inputs are used, conditioned on the output being a failure. Then $3/2=2\bar{p}_{i}+(1+\pi_{i})(1-\bar{p}_{i})$, from which
@@
Thus, conditioned on $\bar{M}_{i}$, the average number of required inputs is
@@
The term $\operatorname{Var}\left[M_{i}\mid\bar{M}_{i}\right]$ in (179) equals the variance of a binomial random variable with parameters $\bar{M}_{i}-r_{i}$ and $\pi_{i}$. Therefore, computing $\operatorname{E}[\bar{M}_{i}]$ from (11) and substituting (180),
@@
$\operatorname{Var}\left[M_{i}\right]=\frac{(14\bar{p}_{i}^{2}-23\bar{p}_{i}+9)r_{i}}{4\bar{p}_{i}^{2}(1-\bar{p}_{i})}=\frac{(9-14\bar{p}_{i})r_{i}}{4\bar{p}_{i}^{2}}.$ (184)
Thus, for $\phi$ small, which implies $\bar{\phi}$ small, substituting (184) into (178) and using (142),
$\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{9r_{1}}{4}\left(\frac{1-14\bar{p}_{1}/9}{\lambda\bar{\theta}}+(1-14\bar{p}_{2}/9)\lambda\bar{\theta}\right)\approx\frac{9r_{1}}{4}\left(\frac{1}{\lambda\bar{\theta}}+\lambda\bar{\theta}\right).$ (185)
The second summand in (177) is computed next. Using the fact that $N_{1}^{\prime}$, $N_{1}^{\prime\prime}$, $N_{2}^{\prime}$, $N_{2}^{\prime\prime}$ are conditionally independent given $s_{1}$, $s_{2}$, taking into account (13) and approximating for $\phi$ small,
@@
$\begin{split}&\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-p_{1})\bar{\phi}^{2}}{\lambda p_{1}^{2}}+\frac{\operatorname{E}[s_{1}-\alpha]p_{1}\bar{\phi}^{2}}{\lambda(1-p_{1})^{2}}+\frac{\lambda\operatorname{E}[s_{2}-\alpha](1-p_{2})\bar{\phi}^{2}}{p_{2}^{2}}+\frac{\lambda\operatorname{E}[s_{2}]p_{2}\bar{\phi}^{2}}{(1-p_{2})^{2}}\\ &\ =\left(\frac{\operatorname{E}[s_{1}]((1-p_{1})^{3}+p_{1}^{3})-\alpha p_{1}^{3}}{\lambda\bar{p}_{1}^{2}}+\frac{\lambda\left(\operatorname{E}[s_{2}]((1-p_{2})^{3}+p_{2}^{3})-\alpha(1-p_{2})^{3}\right)}{\bar{p}_{2}^{2}}\right)\bar{\phi}^{2}\\ &\ =\frac{\operatorname{E}[s_{1}](1-3\bar{p}_{1})-\alpha p_{1}^{3}}{\lambda\bar{\theta}}+\lambda\bar{\theta}\left(\operatorname{E}[s_{2}](1-3\bar{p}_{2})-\alpha(1-p_{2})^{3}\right)\\ &\ \approx\frac{\operatorname{E}[s_{1}]}{\lambda\bar{\theta}}+\lambda\bar{\theta}(\operatorname{E}[s_{2}]-\alpha)\end{split}$ (186)
with $\alpha$ equal to $2$ for OR and 0 for LOR. Using (37) and (38), including the rounding term $\xi$, and then substituting (39)–(42), (135), (136) and (141)–(144), (186) becomes
@@
$\begin{split}\operatorname{E}\left[\operatorname{Var}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}&\approx\frac{1}{\lambda\bar{\theta}}\left(\frac{1}{A}+\mu_{1}+\xi\right)+\lambda\bar{\theta}\left(\frac{1}{A}+\mu_{1}+\xi-\alpha\right)\\ &\quad+2\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right).\end{split}$ (187)
Regarding the third summand, from (11) it stems that $\operatorname{E}[N_{1}]=\operatorname{E}[N_{1}^{\prime}]+\operatorname{E}[N_{1}^{\prime\prime}]=s_{1}/\bar{p}_{1}-\alpha/(1-p_{1})$, and similarly $\operatorname{E}[N_{2}]=s_{2}/\bar{p}_{2}-\alpha/p_{2}$. Then, using (37) and (38),
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\frac{b_{1}^{2}\operatorname{Var}[X]}{\lambda\bar{\theta}}+\lambda\bar{\theta}b_{2}^{2}\operatorname{Var}\left[\frac{1}{X}\right]+2b_{1}b_{2}\operatorname{Cov}\left[X,-\frac{1}{X}\right].$ (188)
For $\phi$ small, $\operatorname{Var}[X]$, $\operatorname{Var}[1/X]$ and $\operatorname{Cov}[X,-1/X]$ are approximately given by the same expressions (162), (163) and (170) as in RR and LRR except with $\theta$ replaced by $\bar{\theta}$ and with $r_{1}=r_{2}$. According to this, and making use of (141)–(144), (40) and (42), (188) yields
$\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\bar{\phi}^{2}l_{1}l_{2}\\ \approx\left(\frac{1}{A}+\frac{3r_{1}}{2}+\mu_{1}+\xi\right)^{\!2}(2r_{1}-1)\cdot\left(\frac{1}{r_{1}(r_{1}-2)}\left(\lambda\bar{\theta}+\frac{1}{\lambda\bar{\theta}}\right)+\frac{2}{r_{1}^{2}}\right).$ (189)
The fourth summand in (177) is bounded using the Cauchy-Schwarz inequality as in Appendix B.1, together with (181) and the fact that $(3-4\bar{p}_{i})/(2(1-\bar{p}_{i}))$ is less than $3/2$ (and approaches that value for $\phi$ small):
$\begin{split}&\left|2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ &\quad\leq 2\sqrt{\operatorname{Var}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}\\ &\quad<3\sqrt{\operatorname{Var}\left[\frac{\bar{M}_{1}}{l_{1}}-\frac{\bar{M}_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.\end{split}$ (190)
The variables $\bar{M}_{1}$ and $\bar{M}_{2}$ are independent. Combined with (13) and (142), this yields
$\operatorname{Var}\left[\frac{\bar{M}_{1}}{l_{1}}-\frac{\bar{M}_{2}}{l_{2}}\right]\bar{\phi}^{2}l_{1}l_{2}=\frac{r_{1}(1-\bar{p}_{1})}{\lambda\bar{\theta}}+r_{2}(1-\bar{p}_{2})\lambda\bar{\theta}\approx r_{1}\left(\frac{1}{\lambda\bar{\theta}}+\lambda\bar{\theta}\right).$ (191)
From (185), (LABEL:eq:_Cov_vasaf_vasas_OR) and (191) it follows that the fourth summand is asymptotically bounded for $\phi$ small as
$\left|2\operatorname{Cov}\left[\operatorname{E}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}\bar{M}_{1},\bar{M}_{2}\right]\!\!,\,\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]\right|\\ \quad<2\sqrt{\operatorname{Var}\left[\frac{M_{1}}{l_{1}}-\frac{M_{2}}{l_{2}}\right]\operatorname{Var}\left[\operatorname{E}\left[\frac{N_{1}}{l_{1}}-\frac{N_{2}}{l_{2}}\mathrel{}\middle|\mathrel{}s_{1},s_{2}\right]\right]}.$ (192)
Using (185), (187), (189) and (192), a similar assessment as in Appendix B.1 can be made: for $\phi$ small; typical values of $A$, for which $1/A$ is large and $r_{1}$ is considerably smaller, namely $r_{1}\approx\sqrt{2/(3A)}$; both for OR ($\mu_{1}=\mu_{2}=2$, $\alpha=2$) and LOR ($\mu_{1}=\mu_{2}=5/4$, $\alpha=0$); and for any $\lambda$ and $\bar{\theta}$, the third summand in the right-hand side of (177) dominates the result for $\operatorname{Var}[\Delta]$. Specifically, the first summand divided by the third takes the following approximate values, where the first corresponds to the case $\lambda\bar{\theta}\approx 1$ and the second to $\lambda\bar{\theta}\gg 1$ and to $\lambda\bar{\theta}\ll 1$:
$\frac{(3r_{1}/2)^{2}}{4(1/A+3r_{1}/2)^{2}},\quad\frac{(3r_{1}/2)^{2}}{2(1/A+3r_{1}/2)^{2}}.$
The ratio of the second and third summands has, for the three cases, the upper bound
$\frac{r_{1}}{2(1/A+3r_{1}/2)}.$
Lastly, the absolute value of the fourth summand in (177) divided by the third is, in view of (192), less than twice the square root of the ratio of first to third summands. Thus, all these ratios are seen to be significantly less than $1$.
It is concluded from the above analysis that the variability of $\Delta$ in OR and LOR, as in RR and LRR, is mostly caused by that of $s_{1}$ and $s_{2}$. This implies that to compute $\operatorname{E}[|\Delta|]$ each variable $M_{i}$, $i=1,2$ in (92) can be replaced by its mean, given by (132), and $N_{i}$ can be replaced by its conditional mean given $s_{i}$, which is approximately $s_{i}/\bar{p}_{i}$ according to (120) and (121). This establishes (152).